# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v2

@@ Abstract
Deep-learning-based techniques have been widely adopted for autonomous driving software stacks for mass production in recent years, focusing primarily on perception modules, with some work extending this method to prediction modules. However, the downstream planning and control modules are still designed with hefty handcrafted rules, dominated by optimization-based methods such as quadratic programming or model predictive control. This results in a performance bottleneck for autonomous driving systems in that corner cases simply cannot be solved by enumerating hand-crafted rules. We present a deep-learning-based approach that brings prediction, decision, and planning modules together with the attempt to overcome the rule-based methods’ deficiency in real-world applications of autonomous driving, especially for urban scenes. The DNN model we proposed is solely trained with 10 hours of valid human driver data, and it supports all mass-production ADAS features available on the market to date. This method is deployed onto a Jiyue test car with no modification to its factory-ready sensor set and compute platform. The feasibility, usability, and commercial potential are demonstrated in this article.  [^1] Test records will be available at https://github.com/weijian-sun/GAD
@@ 1 Introduction
Autonomous driving software on the market to date is usually designed following a classical engineering methodology called “divide and conquer”, where the problem is divided into smaller sub-tasks to reduce complexity. These sub-tasks usually include perception, prediction, motion planning, and control. However, this easily interpretable rule-based method is highly dependent on accurate prior knowledge, which severely restricts its scalability. High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information, greatly lessens the burden of perception and enormously facilitates planning tasks for autonomous driving. Unfortunately, HD maps are proven hard to scale up due to the complexity and costliness of map generation and maintenance, as well as failure to reflect changes in real-world road conditions timely, which is often constrained by resource-intensive collection processes. Furthermore, the rule-based planning module is held accountable for most failures and deficiencies of the autonomous driving system in real driving scenarios, especially for urban scenes. Hence, it is often regarded as the bottleneck of the entire software stack. New rules must be designed and re-tuned in accordance with different driving scenarios and new corner cases. This process is expensive and it scales poorly for the simple fact that the scenario space for the Autonomous Drive Vehicle(ADV) is infinite while rules can only be of limited number.
@@
End-to-end autonomous driving, which aims to consider sub-tasks in a full-stack manner jointly, has been proposed to resolve the above issues. Methods[22, 24, 25, 40, 20, 33, 5, 30] that direct learning from raw sensor data to planning trajectories or driving commands eliminate the need for hand-crafted heuristic rules and the prerequisite HD maps, avoiding infomation loss and better adapt to diverse scenarios. However, these approaches lack interpretability which plays a key role in understanding and tracing how the system envolves, thus will be difficult for the developer to optimize. Interpretability is of key importance in a safety-critical system, particularly if a bad event were to happen in mass production. Moreover, the absence of modularized structure and prior knowledge makes these methods very brittle when deployed in real-world closed-loop.
@@
To overcome the aforementioned problems, we explore an alternative method that runs entirely on the edge and purely relies on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [12] based on perception results. The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first, primarily due to considerations of scalability and mass production feasibility. We believe this is effective for two primary reasons. Firstly, landmarks provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized. Secondly, for turning scenes, we gather routing information from historical commuting trajectories and avoid collisions with the help of real-time occupancy. However, the proposed methodology is not limited to commuting scenarios; it supports a versatile selection of route formats. We also validated its non-commuting variant in the experiments.
@@
To break the bottleneck brought by rule-based planning but maintain the system’s explainability and stability, we propose a general framework to handle the prediction-planning tasks in a data-driven and integrated manner. Our framework consists of a trajectory generator and evaluator. The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric info in grid form to validate and score candidates [^2] We note that by parametric form we mean the structured numerical representations of certain perception outputs, such as the observed vehicle states. Conversely, the non-parametric form refers to unstructured representations that cannot be described by parameters, such as occupancy.. We validate our framework in a closed loop by deploying and testing it in real-world user-level autonomous driving vehicles.
We elucidate the rationale behind our architectural that segregates perception from the comprehensive end-to-end stack: Firstly, for mass production friendly concern, the modularized perception provides explaninablity and tunablitiy, Secondly, information articulated in parametric form facilitates reasoning at the agent level, not solely for application within the safety layer and the analysis of explainable cases, but also to enhance our trajectory generation processes. Thirdly, parameterization encompasses post-tracking and fusion for the results of single or multi-modal detection on raw sensor data which remain formidable unresolved issues, particularly when considering the substitution of these processes with a model to attain comparable results. We leave perception involved version to follow up works when those issue resolved in the future.
The real-world closed-loop validation needs to be highlighted as well: Firstly, a on road planer will adjust to the target in the loop to keep safe and comfortable since real-world planning needs handle the deviations brought by the external disturbances and controller’s modeling errors, which is hard to simuate and omitted by open-loop planner as well. Secondly, the validity of the trajectory cannot be verified from the perspective of the downstream controller. Since unexpected and unreasonable jerks and curvatures can only be tested when the controller is in the loop with the planner. The above limitations make the results in open-loop or simulator less convincing and leaves the effectiveness of those methods (like UniAD[22], VAD[24], mile[20], ST-P3[21]) unknown for onboard road tests and mass production.
Our key contributions are summarized as follows:
  1. 1.
    We propose an industry-level paradigm for data-driven prediction-planning with an HD map-free setting. Our model leverages non-parametric information with better tolerance to upstream perception and handling prediction and planning tasks in one backbone.
  2. 2.
    To the best of our knowledge, we are the first evaluation of such system in a complex, real-world, urban road with a factory-ready sensor set and compute platform. We demonstrate the imperative of closed-loop evaluation compared to imitation metric adopted by extensive offline studies.
  3. 3.
    We extend the state-of-the-art data-driven approach and propose a method that goes beyond simple imitation or trajectory sampling and instead reasons about the potential in generative learning by combining max marge planning with multi-modal imitation. As depicted in Fig. 1
@@ 2.1 End to End Learning&Validation
End-to-end autonomous driving has become a significant research focus, with numerous recent studies. These methods, using neural networks, produce planning trajectories or control commands, with many[40, 47, 20] focusing on closed-loop simulator experiments [10, 27]. However, a persistent domain discrepancy exists between the simulated environment and the actual world, especially in terms of sensor data fidelity and the dynamic states of agents. The accurate modeling of agent behaviors and the simulation of real-world data[14] within closed-loop environments continue to pose formidable challenges that have yet to be resolved. Recently, open-loop experiments with real-world data have attracted more attention[24, 22, 21, 48]. Methods like UniAD[22] claim their effectiveness in improving final planning performance by involving learning intermediate tasks. However, it relies on an additional optimal control module to obtain a feasible trajectory and prevent collisions, which violates the initial purpose of breaking optimization-based deficiencies in motion planning task. Nevertheless, the effectiveness of open-loop demonstration has been challenged by [28] in terms of imbalanced data distribution and metric limitation. We address these contentious issues by rigorously validating our approach through real-world closed-loop road test, wherein the proposed methodology eschews the dependence on conventional optimization algorithms to derive trajectories that adhere to specified constraints.
@@ 2.2 Data Driven Prediction
Prior research has extensively explored actor-based prediction methods for future driving dynamics. These methods either produce specific trajectories [4, 2, 32, 51] or use generative models to outline probable outcomes [36, 41, 17, 26]. In order to encode scene context, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8]. More recent methods have started to leverage the topologic graph obtained from HD maps in order to better represent lane connectivity. VectorNet [13] encodes map and agent as polylines with a global interaction graph. LaneGCN [29] treats actor past and the lane graph separately and then fuses them with a series of attention layers. Temporal occupancy maps [23, 37, 19], on the other hand, provide another option to forecast grid or heat map at the scene level. We follow the philosophy of [11, 18, 38, 42, 46] in modeling scene-level representations but propose a multi-stage pipeline to perform forecasting in both formats. Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are mainly adopted in the safety layer and to boost some of our samplers.
@@ 2.3 Data Driven Planning
For imitation based methods, modeling strategies from trajectory prediction are adopted, which build an imitation decoder trained under the supervision of human demonstration. Apart from the aformention end to end works, SafetyNet[44] developed a machine learning planner and tested it in downtown San Francisco, but the model can hardly avoid collision without the help of a fallback layer. VAD[24] modeled the driving scene as a fully vectorized representation. PlanT[35] proposed a transformer-based backbone in actor-level representation and validated in simulation. However, the monotonous single trajectory imitation suffers from robustness and lacks interpretability[6], thus difficult to tune by case and hard to scale up. More recently, explicit methods with great interpretability have been shown to adapt better to more challenging environments. Such approaches usually build a cost volume with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost. The cost volume is organized in a non-parametric form learned directly from the network[48, 38, 3, 49, 7]. Beginning with pioneering work on NMP [48], DSDNet[49] and P3[38] made further developments that combined hand-made and learning-based costs to obtain an integrated cost volume. In the follow-up work, MP3 [3] introduced a mapless approach, while ST-P3[21] explored vision-based inputs. We also adopt this combination to choose the best trajectory but extend the system’s robustness by boosting the generating algorithm with modularized perception output. In contrast to all methods above, which solely rely on a rule-based sampler, our model leverages multi-modal imitation learning to build a diversified hybrid trajectory set and stays safe by being evaluated in an interpretable cost volume as a summary of scene occupancy and agent motion.
Figure 2: Model Structure of GAD. The model initiates with two streams for rasterization and vectorization respectively to encode scene and instance-level features. The rasterization branch utilizes convolution layers to derive the scene embedding, which is then decoded into cost volumes for scoring candidate trajectories from multiple samplers and prediction grid maps for agent prediction. Meanwhile the scene embedding are cropped and incorporated with the instance encoding from vectorization branch to decode ego trajectory and agents’ future movements.
@@ 3.1 Input Representation
The goal of this paper is to present a deep learning (DL) pipeline that goes from prediction to motion planning in an integrated manner; hence, it is not our intention to cover perception at length. In our work, the perception module follows a typical mainstream design that provides tracked entities, landmarks, and occupancy voxels (also known as volumn pixels) by incorporating images from a suite of independent cameras with different field of views (FoVs) to formulate a surround view at 360 degrees. The perception module has a deep neural net that directly stitches features in bird’s-eye view (BEV) form, and transformers are used to capture both spatial and temporal contexts. Occupancy is also an integral part of outputs by this perception module, providing a useful scene representation of a three-dimensional world that offers free space for SDV to navigate through. More details are provided in the appendix. To encode upstream output for different usages, we design representations for scene and instance levels [^3] We note that by scene we mean the overall environment context for the current frame, by instance we mean the elements in the scene such as ego, agent, landmark, route, etc., which are applied to trajectory evaluation and generation, respectively.
@@
We adopt rasterization proposed by[9] for scene representation. Scene rasterization with convolutional forwarding can better capture scene-level cues compared to instance vectorization, avoiding information loss to the maximum extent. Owing to the convolutional network’s property of exploiting spatially local correlation, undefined or irregular elements like construction areas, traffic cones, and roadblocks can be incorporated with a unified design regardless of how they are organized and whether they have regularized shapes. Additional scene information depicted by free space or occupancy, lacking parametric description, can hardly be involved with vectorized representation and, once omitted, can be fatal to safety-critical systems. To serve as a qualified trajectory evaluator, BEV rasterization involves environment cues from online perception outputs as much as possible to capture those “devils in detail”. We follow the interpretable design as that of in [3], but rasterize different scene elements(e.g. agent, lanes, occupancy, route) into different binary channels to enable reasoning about the distinct element separately. The details can be found in our appendix.
@@
The instance representation is employed to address the shortcomings of scene encoding, which are afflicted by the truth that, rasterization is prone to distorting highly organized and structured information, rendering it disorganized. Since this procedure is irreversible, making it challenging to retrieve instance encoding for lane center lines, road boundaries, and other elements. The vectorized representation combined with the transformer has been proven to be superior to CNN forwarding on BEV raster in terms of accuracy and efficiency[13] for the motion prediction task. To take advantage of modeling in such form, we adopt vectorization[13] to encode the perception outputs into vector sets. Instances with parametric description are encoded in the vector form, including route points, ego and agent historical, lane center lines, and road boundaries, while non-parametric outputs are ignored, like occupancy, construction area, roadblocks, and other irregular elements. Each element includes position and type features, with the position described relative to the ADV pose. The relationships between separated elements are organized as a graph.
@@ 3.2 Model Structure
We build a two-stream backbone network to encode inputs and extract agents and scene features separately. As shown in Fig. 2. One steam processes the rasterized scene while the other deals with the vectorized graph. To elaborate, the rasterization stream employs a convolutional encoder of multiple resolutions to derive the scene embedding, responsible for aggregating geometric information from the scene context to extract global appearance and motion cues. The scene embedding is then passed through two separate deconvolution decoders to generate the cost volumes and the predicted grid maps respectively. Meanwhile, in the other stream, inspired by [13], the instance-level encoding is built on a hierarchical graph network-based architecture. It consists of a PointNet-based [34] local sub-graph for processing local information from vectorized inputs and a global graph using a Transformer encoder [43] for reasoning about interactions over agents and landmark features. Finally, the corresponding scene encoding for the agent is derived via oriented crop[4] in ego-centric scene feature and fused with instance encoding to obtain agent-level features $f_{agent}$. Then an imitation decoder is utilized to produce the trajectories for agents and the ego. The details for the encoders and decoders are illustrated in the appendix.
@@ 3.3 Route
When HD maps are accessible, the input route is typically given in the form of a sequence of lanes that the ADV should follow. However, in HD map-free settings, this is not possible, especially for turning scenarios in intersections where no leading landmark is available. In our mapless setting, route information is derived from visual landmark and human commuting trajectories for different scenarios to substitute those absent cues. The lane center lines constructed by landmarks are mainly used in cruise and lane change scenarios, while in turning scenarios, the starting and ending positions are recorded in advance, once vehicles approach the intersection, the human commuting trajectories will serve as route. More details can be found in the appendix.

### v3

@@ Abstract
Deep-learning-based techniques have been widely adopted for autonomous driving software stacks for mass production in recent years, focusing primarily on perception modules, with some work extending this method to prediction modules. However, the downstream planning and control modules are still designed with hefty hand-crafted rules, dominated by optimization-based methods such as quadratic programming or model predictive control. This results in a performance bottleneck for autonomous driving systems in that corner cases simply cannot be solved by enumerating hand-crafted rules. We present a deep-learning-based approach that brings prediction, decision, and planning modules together with the attempt to overcome the rule-based methods’ deficiency in real-world applications of autonomous driving, especially for urban scenes. The DNN model we proposed is solely trained with 10 hours of valid human driver data and supports all mass-production ADAS features available on the market to date. This method is deployed onto a Jiyue test car without modification to its factory-ready sensor set and compute platform. The feasibility, usability, and commercial potential are demonstrated in this article.
@@ 1 Introduction
Autonomous driving software on the market to date is usually designed following a classical engineering methodology called “divide and conquer”, where the problem is divided into smaller sub-tasks to reduce complexity. These sub-tasks usually include perception, prediction, motion planning, and control. However, this easily interpretable rule-based method is highly dependent on accurate prior knowledge, which severely restricts its scalability. High-definition (HD) maps, among various forms of prior knowledge, play a crucial part in offering rich semantic information. Unfortunately, HD maps are proven hard to scale up due to the complexity and costliness of map generation and maintenance, as well as failure to reflect changes in real-world road conditions timely, which is often constrained by resource-intensive collection processes. Furthermore, the rule-based planning module is held accountable for most failures and deficiencies of the autonomous driving system in real driving scenarios, especially for urban scenes. Hence, it is often regarded as the bottleneck of the entire software stack. New rules must be designed and re-tuned in accordance with different driving scenarios and new corner cases. This process is expensive and it scales poorly for the simple fact that the scenario space for the Autonomous Driving Vehicle(ADV) is infinite while rules can only be of a limited number.
@@
End-to-end autonomous driving, which aims to consider sub-tasks in a full-stack manner jointly, has been advocated to address the aforementioned challenges. Methods[23, 25, 26, 43, 21, 35, 5, 32] that facilitate direct learning from sensor data to trajectories or driving commands obviate the necessity for manually devised heuristic rules and the dependency on HD maps. However, these methodologies are deficient in interpretability, which is crucial for elucidating and monitoring the system’s evolution, consequently complicating optimization efforts by developers. Moreover, the absence of modularized structure and prior knowledge makes these methods exceedingly brittle when operationalized in real-world closed-loop.
@@
To overcome the aforementioned problems, we explore an alternative method that runs entirely on the edge and purely relies on on-the-fly computation of local topology reconstructed with landmarks and occupancy grids [13] based on perception results. The reconstruction is designed to support motion planning in urban scenarios, and we explore its practicality for daily commuting first, primarily due to considerations of scalability and mass production feasibility. However, the proposed methodology is not limited to commuting scenarios; it supports a versatile selection of route formats. We also validated its non-commuting variant in the experiments.
@@
To break the bottleneck brought by rule-based planning, we propose an interpretable and tunable framework to handle the prediction-planning tasks in a data-driven and integrated manner. Our framework consists of a trajectory generator and evaluator. The generator utilizes vectorized embedding from parametric input to perform trajectory generation, while the evaluator leverages rasterized non-parametric information in grid form to validate and score candidates [^1] We note that by parametric form we mean the structured numerical representations of certain perception outputs, such as the observed vehicle states. Conversely, the non-parametric form refers to unstructured representations that cannot be described by parameters, such as occupancy.. We validate our framework in a closed loop by deploying and testing it in real-world user-level autonomous driving vehicles.
Our key contributions are summarized as follows: (1) We introduce a scalable, tuneable, industry-grade framework for data-driven prediction-planning in HD map-free setting. Our model offers an alternative to the e2e model and demonstrates performance on par with off-the-shelf, optimization-based planner. (2) To the best of our knowledge, we are the first evaluation of such system in a complex, real-world, urban road with factory ready sensor set and compute platform [^2] Road test video is available at https://youtu.be/DYPB5lCRvn0. We demonstrate the imperative of closed-loop evaluation compared to imitation metric adopted by extensive offline studies. (3) We extend the state-of-the-art data-driven approach by proposing a method that goes beyond simple imitation or trajectory sampling. Instead, our method reasons about the potential in generative learning by combining max margin planning with multi-modal imitation, as depicted in Fig. 1
@@ 2.1 End to End Learning&Validation
End-to-end autonomous driving has become a significant research focus. These methods, using neural networks, produce planning trajectories or control commands, with many[43, 50, 21] validations on closed-loop simulator[11, 29]. However, a persistent domain discrepancy exists between the simulated environment and the actual world, the accurate modeling of agent behaviors and the simulation of real-world data[15] within closed-loop environments continue to pose formidable challenges that have yet to be resolved. Recently, open-loop experiments with real-world data have attracted more attention[25, 23, 22, 51]. Methods like UniAD[23] claim their effectiveness in improving final planning performance by involving learning intermediate tasks. Nevertheless, it necessitates an auxiliary optimization-based module to generate a viable trajectory and avert collisions, thereby contravening the foundational objective of circumventing deficiencies inherent in optimization-based approaches. Moreover, the effectiveness of open-loop demonstration has been challenged by [30] in terms of imbalanced data distribution and metric limitation. Those unresolved limitations make the results in open-loop or closed-loop simulator less convincing and leaves the effectiveness of methods (like UniAD[23], VAD[25], mile[21], ST-P3[22]) unknown for onboard road tests and mass production. We address these contentious issues by rigorously validating our approach through real-world closed-loop road test, wherein the proposed methodology eschews the dependence on conventional optimization-based algorithms.
@@ 2.2 Data Driven Prediction
Prior researches have extensively explored actor-based prediction methods for future driving dynamics[4, 2, 34, 54, 38, 44, 18, 28]. In order to encode scene contexts, a classic CNN architecture can be applied to a rasterized map to predict 2D coordinates [8]. More recent methods have started to leverage the topologic graph obtained from HD maps in order to better represent lane connectivity. VectorNet [14] encodes map and agent as polylines with a global interaction graph. LaneGCN [31] treats actor histories and the lane graph separately and then fuses them with a series of attention layers. Temporal occupancy maps [24, 39, 20], on the other hand, provide another option to forecast grid or heat map at the scene level. We follow the philosophy of [12, 19, 41, 45, 49] in modeling scene-level representations but propose a multi-stage pipeline to perform forecasting in both formats. Moreover, unlike typical methods, which heavily rely on the precision of trajectory output, our agent results are primarily used to enhance the safety layer and to boost some of our samplers.
@@ 2.3 Data Driven Planning
Apart from the aformention end to end works, imitation-based methods have been widely explored in the planning task. SafetyNet[47] developed a machine learning planner and tested it in downtown San Francisco, but failed to avoid collision without the help of a fallback layer. VAD[25] modeled the driving scene as a fully vectorized representation. PlanT[37] proposed a transformer-based backbone in actor-level representation and validated in simulation. However, the monotonous trajectory regression suffers from robustness and lacks interpretability[6], thus difficult to tune by case and hard to scale up. More recently cost-volume-based methods have improved adaptability and interpretability in challenging environments. They identify optimal trajectories by evaluating candidates within a network-learned cost volume[51, 41, 3, 52, 7], while the candidate trajectories are generated through rule-based samplers. Beginning with pioneering work on NMP [51], DSDNet[52] and P3[41] made further developments that combined hand-made and learning-based costs. In the follow-up works, MP3 [3] introduced a mapless approach, while ST-P3[22] explored vision-based inputs. We also adopt this combination to choose the best trajectory, but in contrast to all methods above, which solely rely on a rule-based sampler, we explore the potential of multi-modal imitation and generative learning in motion planning task.
Figure 2: The Model Structure of GAD initiates with two streams, rasterization and vectorization, to encode scene and instance-level features; the rasterization branch uses convolution layers to derive the scene embedding, which is decoded into cost volumes and prediction grid maps, while the scene embedding is cropped and combined with the instance encoding from the vectorization branch to form $f_{agent}$ which decoded into ego trajectory and agents’ future movements.
@@ 3.1 Input Representation
The goal of this paper is to present a deep learning (DL) pipeline that goes from prediction to motion planning in an integrated manner; hence, it is not our intention to cover perception at length. In our work, the perception module follows a typical mainstream design that provides tracked entities, landmarks, and occupancy voxels (also known as volumn pixels) by incorporating images from a suite of independent cameras with different field of views (FoVs) to formulate a surround view at 360 degrees. More details are provided in Sec. A.1. To encode the upstream output for diverse applications, we have devised representations at both the scene and instance levels [^3] Here, by scene, we refer to the comprehensive environmental context pertinent to the current frame, and by instance, we denote the constituents within the scene such as the ego vehicle, other agents, landmarks, routes, etc., which are systematically utilized in the evaluation and generation of trajectories.
@@
We employ the rasterization methodology as proposed by [10] for the representation of scenes. Rasterization, when coupled with convolutional forwarding, more effectively captures scene-level cues than instance vectorization. Due to the inherent ability of convolutional networks to exploit spatial correlations, it is feasible to integrate undefined or irregular elements such as construction zones, traffic cones, and roadblocks into a cohesive design, irrespective of their arrangement or geometric regularity. Furthermore, scene information characterized by free space or occupancy, which lacks a parametric framework, is challenging to incorporate within a vectorized representation and, if excluded, poses significant risks to safety-critical systems. BEV rasterization assimilates environmental cues from online perception outputs to the fullest extent to capture those ’devils in detail’. Comprehensive details are provided in Sec. A.2.
@@
The instance representation is utilized to mitigate the deficiencies inherent in scene encoding, which suffers from the fact that rasterization tends to distort highly organized and structured data into non-parametric forms. Given that this process is irreversible, it complicates the extraction of instance encoding for elements such as lane centerlines, road boundaries, among scene feature map. On the other hand, vectorized representation combined with the transformer has been proven to be superior to CNN forwarding on BEV raster for instance reasoning[14]. To leverage the modeling capabilities afforded by this form, instances with parametric description are encoded in the vector form, including route points, agent histories, lane center lines, and road boundaries, while non-parametric outputs are ignored, such as occupancy, construction area, roadblocks, and irregular elements. More details can be found in Sec. A.3
@@ 3.2 Model Structure
We build a two-stream backbone network to encode inputs and extract agents and scene features separately. As shown in Fig. 2. One steam processes the rasterized scene while the other deals with the vectorized graph. To elaborate, the rasterization stream employs a convolutional encoder of multiple resolutions to derive the scene embedding, which is then passed through two separate deconvolution decoders to generate the cost volumes and the predicted grid maps respectively. Meanwhile, in the other stream, inspired by [14], the instance-level encoding is built on a hierarchical graph network-based architecture. It consists of a PointNet-based [36] local sub-graph for processing local information from vectorized inputs and a global graph using a Transformer encoder [46] for reasoning about interactions over agents and landmark features. Finally, the corresponding scene encoding for the agent is derived via oriented crop[4] in ego-centric scene feature and fused with instance encoding to obtain agent-level features $f_{agent}$. Then an imitation decoder is utilized to produce the trajectories for agents and the ego. The details structure for the encoders and decoders are illustrated in Sec. A.5.
@@ 3.3 Route
When HD maps are available, the route is typically a sequence of lanes for the ADV to follow. In settings without HD maps, this approach is not feasible. In this work, route information is derived from visual landmarks and human commuting trajectories for different scenarios to substitute those absent cues. This approach is deemed effective for two primary reasons. Firstly, landmarks (lane center, lane divider and road boundary) provide sufficient cues for driving in a cruise scenario - it generally consists of lane keep and lane change for following and overtaking behaviors regardless of how road topologies are organized. Secondly, for turning scenes(left turn, right turn, u-turn), we gather routing information from historical commuting trajectories, supplemented by real-time occupancy data to mitigate collision risks. The implementation details can be found in Sec. A.6.